# Ctrl-V: Higher Fidelity Video Generation with Bounding-Box Controlled Object Motion

### v2

@@ 5 Conclusions
We present Ctrl-V, a novel model capable of generating controllable autonomous vehicle videos. Our approach demonstrates that the BBox Predictor model can closely follow generation requirements for the first and last frames and produce a coherent bounding-box track for the intermediate frames. We also show that our Box2Video  network generates high fidelity videos, strictly adhering to the given bounding boxes. Furthermore, our model accommodates both 2D and 3D bounding boxes and handles uninitialized objects appearing in the middle of the videos. Ctrl-V provides future researchers with an efficient way to simulate driving video data with flexible controllability in the form of bounding boxes. In addition, we further define an improved metric to evaluate bounding box conditioned video generation to account for objects that are not present in the first frame, and those that do not remain until the last frame. In Appendix D, we discuss potential future work for this project. With Ctrl-V and an improved metric for more accurate evaluation, we aim to establish a solid foundation for future research in controllable video generation.
Acknowledgments
We thank CIFAR for their support under the AI chairs program, NSERC for their support under the Discovery Grants program, and Samsung for supporting this work.
Appendix A Implementation Details
A.1 Step-by-Step Guide to Plotting Bounding Box Frames
We create bounding-box plots for each frame according to the following steps:
@@ A.1 Step-by-Step Guide to Plotting Bounding Box Frames
    For each object, we fill the 2D bounding-box region with its track ID’s color at 75% transparency. This transparency allows the overlapping regions of the boxes to remain visible.
@@
    If 3D bounding-box information is available for an object, we draw the 3D bounding boxes and outline them with the color corresponding to the object’s class label.
@@
    If 3D bounding-box information is unavailable, we outline the 2D bounding-box with the color corresponding to the object’s class label. Moreover, we mark an X at the rear face of each 3D bounding-box for enhanced contextualization.
@@
    In certain experiments, we substitute bounding-box plots with trajectory plots. To create a trajectory plot, we first compute the mid-points of its 2D bounding boxes. Subsequently, at each bounding-box midpoint, we draw a circle with a diameter of 10 pixels, filled with the color corresponding to its track ID. Within this circle, we draw a smaller circle with a diameter of 5 pixels, filled with the color representing the object’s class label.
A.2 Bounding-Box Visualization and Reconstruction Analysis
In this work, we refrain from introducing new modules dedicated to encoding the bounding-box frames. Instead, we leverage the encoder ($\mathcal{E}$) within the SVD’s autoencoder framework. Our experiments demonstrate the efficacy of the SVD autoencoder in accurately encoding and decoding bounding-box frames.
@@ A.2 Bounding-Box Visualization and Reconstruction Analysis
Figure 4: Visualizing a BDD100K 2D-bbox frame created using the method described in A.1
@@
Figure 5: Visualizing a VKITTI 3D-bbox frame created using the method described in A.1
Figure 6: Visualizing a BDD100K trajectory frame created using the method described in A.1.
A.3 Dataset Configuration
KITTI: The downloadable content of the official KITTI dataset does not include bounding-box label files for the testing set. As a result, we manually partition the dataset’s training folder into two segments for our experiments. The training split encompasses samples ‘$0$’ to ‘0018’, while the testing split consists of samples ‘0019’ and ‘0020’.
@@ A.3 Dataset Configuration
In our experiments, we standardize the frame sizes for all experiments to $312\times 520$ and set a clip length of 25 frames. All experiments are trained using the training set and assessed using the validation or test sets. Unless otherwise stated, the results and the visualizations provided are derived from the test sets.
A.4 Training Configuration
Our pretrained SVD models are adaptations from HuggingFace’s diffuser library [45]. Specifically, we utilize the ‘stable-video-diffusion-img2vid-xt’ SVD variant in this project.
@@ A.4 Training Configuration
The 3D bounding-box predictor anticipates frames using 3D bounding boxes, while the 2D predictor does the same for 2D bounding boxes. A 1-to-1 model predicts frames between the first and last bounding-box frames, while a 3-to-1 model predicts frames between the first 3 and last bounding-box frames. Additionally, there’s a bounding-box-trajectory model where the last conditional frame is a trajectory frame instead of a bounding-box frame.
@@
For the bounding-box predictors trained on KITTI or vKITTI datasets, each model undergoes 5 epochs of training. The 2D bounding-box predictor (1-to-1) is trained on BDD100K for 33,000 iterations. Additionally, the 2D bounding-box (3-to-1) and 2D bounding-box-trajectory (3-to-1) models start their fine-tuning process for an additional 15,000 iterations from where the 2D bounding-box predictor (1-to-1)’s training ends. All of these prediction models are trained on a single A100 GPU with a batch size of 1. We also set gradient accumulation steps to 5 and employed AdamW [46] as our optimizer.
@@
The fine-tuning and training procedures for the baseline Stable Diffusion and the ControlNet models are almost the same as of the bounding-box predictor model. The only difference is that we trained our baseline and ControlNet for 48,000 iterations on BDD100K.
Appendix B Evaluation Details
B.1 Generation Quality Metrics: PSNR, SSIM, LPIPS and FVD
@@
PSNR, or Peak signal-to-noise ratio, calculates the ratio between the maximum value of a signal and the noise that affects the fidelity of its representation. It is closely related to Mean Squared Error (MSE) loss. It can be in fact calculated by subtracting the log of MSE from a constant. It is long been used as a metric for image reconstruction quality for its ease of use and mathematical properties. However, PSNR suffers multiple issues as a metric for image generation tasks. Its insensitiveness to various distortion such as blurring, pepper noise, and mean-shifting means that image pairs that look very different to human can have very similar PSNR scores [47, 48].
@@
SSIM, or Structural similarity index measure, is a metric that aims to evaluate similarity in ”structures” of image pairs – as opposed to absolute errors as in the cases of MSE and PSNR. Because the human visual system is more sensitive to structural distortions, this metric is more closely aligned with human evaluation compared to PSNR. However, some visually apparent distortions are not captured by the SSIM. Things like changing hue and brightness do not seem to affect SSIM much [49] and Sharif et al. [50] presents more examples of unpredictable ssim behaviours when presented with different distortions.
@@
LPIPS, or Learned Perceptual Image Patch Similarity [37], is a neural network based approach to accessing image similiarity. It does so by computing distance between some reprensentation of image patches. These representations are obtained by running image patches through some pre-defined neural network. LPIPS has been shown to math human perception well.
@@
FVD, or Frechet Video Distance [36], is a metric for generative video models. It is based on the Fréchet inception distance (FID) for images. Unlike the 3 other metrics that focuses on image frames and do not consider their temporal relationships, FVD is specifically designed for videos. It takes into consideration the visual quality, temporal coherence, and diversity of samples of videos generated by the model under evaluation. The large scale human study carried out by the authors suggest FVD consistently agrees with human evaluation of videos.
B.2 Prediction Scores: Converting Bounding-Box Frames to Binary Masks
@@
To transform predicted bounding-box frames into binary masks, we first need to perform post-processing on the frames. These frames originate from a network output, and although background pixels that seem black, they may not all be precisely zero. To prevent these pixels being detected as bounding-box masks, we aggregate the pixel values across color channels, and for those channels whose sum is below 50, we convert them to black pixels. Subsequently, we convert the frames into binary masks by aggregating all channels and changing the colored pixel values to 1.
@@
Accurately evaluating bounding-box predictions remains challenging. We recognize that our current metrics (maskIoU, maskR and maskP), which rely on binary masks to evaluate the performance of bounding-box predictors, have limitations as they do not consider object tracking IDs (the colors of the bounding boxes). However, we still believe that the metric evaluations using binary masks can offer valuable insights into our model’s performance.
B.3 Motion Control Assessment: YOLOv8 Object Detector Configurations
@@
We utilize the “yolov8x” variant of the YOLOv8 model, implemented by Ultralytics [51], for object detection. We set the detector’s Non-Maximum Suppression Intersection Over Union (NMS-IoU) threshold to be 0.35 across all experiments.
Since YOLOv8 was trained to detect objects from the MS COCO dataset [40], we adjust the class labels of our dataset to match the MS COCO labels using the following mappings:
@@
We have attempted to enhance our detection results across frames by using a tracker in conjunction with the YOLOv8 detector. Specifically, we use the BoT-SORT [52] tracker. However, we’ve observed an increase in the precision but a significant drop in the recall of the detection scores. Thus, we’ve reverted to using only the YOLOv8 detection model to compute the AP metrics.
B.4 Motion Control Assessment: Average Precision Calculation
@@
We begin by employing a YOLOv8 detection model on both the generated and ground-truth videos, yielding a series of bounding-box detections for each. When we generate bounding-box detections on ground-truth videos, we keep the confidence cutoff fixed at 0.6.
Next, we calculate the AP scores at different IoU thresholds following the MS COCO protocol [40]: compute 10 AP scores at IoU thresholds ranging from 0.5 to 0.95, in increments of 0.05. The AP score computed at threshold $\tau$ is referred to as $\text{AP}_{\tau\times 100}$. Mean average precision (mAP) is the average of all the AP scores. The computed results are reported in Table 4.4.
@@
Appendix C Generation Results
C.1 Stable Video Diffusion vs Box2Video: Generation Quality Comparison
The following are generation comparisons between the fine-tuned Stable Video Diffusion baseline model and the Box2Video conditioned on the ground-truth bounding-box frames. All visualizations are generated using the BDD100k dataset.

### v3

@@ 5 Conclusions
We have presented Ctrl-V, a novel model capable of generating controllable autonomous vehicle videos via bounding box trajectory conditioning. Our approach demonstrates that our BBox Generator  technique can closely follow generation requirements for the first and last frames and produce a coherent bounding box track for intermediate frames. Moreover, our Box2Video network generates high-fidelity videos that strictly conform to the provided bounding boxes. Furthermore, our model accommodates both 2D and 3D bounding boxes and handles uninitialized objects appearing in the middle of the videos. Ctrl-V provides future researchers with an efficient way to simulate driving video data with flexible controllability in the form of bounding boxes. In addition, we further define an improved metric to evaluate bounding box conditioned video generation to account for objects that are not present in the first frame, and those that do not remain until the last frame. In Appendix G, we discuss potential future work for this project. With Ctrl-V and an improved metric for more accurate evaluation, we aim to establish a solid foundation for future research in controllable video generation.
Appendix A Video Demonstrations
For a more extensive visual presentation of our findings, including generated videos, please follow this link: https://anongreenelephant.github.io/ctrlv.github.io/.
Appendix B Notations
B.1 SVD Scaling Function Definitions
As presented in Equation 2, $\lambda_{\text{skip}}$, $\lambda_{\text{out}}$, $\lambda_{\text{in}}$ and $\lambda_{\text{noise}}$ are scaling functions where $\lambda_{\text{in}}$ and $\lambda_{\text{out}}$ scale the input and output magnitudes of the neural network ${\mathbb{U}}_{\theta}$ being trained, $\lambda_{\text{skip}}$ modulates the skip connection, and $\lambda_{\text{noise}}$ maps noise level $\sigma_{t}$ into a conditioning input for ${\mathbb{U}}_{\theta}$. More detailed information on these functions can be found in the work of Karras et al. (2022).
Unless otherwise specified, we use the SVD definitions as presented in (Blattmann et al., 2023a). The mathematical formulations for the scaling functions are thus:
$\begin{split}\lambda_{\text{out}}(\sigma)=\frac{-\sigma}{\sqrt{\sigma^{2}+1}}&\quad\quad\lambda_{\text{in}}(\sigma)=\frac{1}{\sqrt{\sigma^{2}+1}}\\ \lambda_{\text{skip}}(\sigma)=(\sigma^{2}+1)^{-1}&\quad\quad\lambda_{\text{noise}}(\sigma)=\frac{\log\sigma}{4}\end{split}$ (3)
B.2 Ctrl-V Notations
In this section, we provide a table of definition for all of the Symbols notations used in our work.
Symbol Appearance Definition
$\mathcal{E}$ Section 3.1, Figure 2 the off-the-shelf VAE Image Encoder used by SVD
$\mathcal{D}$ Section 3.1, Figure 2 the off-the-shelf VAE Image Decoder used by SVD
$\mathcal{C}_{\xi}$ Figure 2 modified ControlNet module in Box2Video
${\mathbb{D}}_{\theta}$ Equation 2 denoiser network of the SVD backbone
${\mathbb{U}}_{\theta}$ Equation 1, Equation 2 the main UNet component of ${\mathbb{D}}$
${\bm{f}}$ Section 3.1 a sequence of frames
${\bm{f}}^{(i)}$ Section 3.1 the $i^{th}$ frame in ${\bm{f}}$
${\bm{z}}$ Section 3.1 ${\bm{z}}=\mathcal{E}({\bm{f}})$. a sequence of frames encoded by the VAE encoder
${\bm{z}}^{(i)}$ Section 3.1 ${\bm{z}}^{(i)}=\mathcal{E}({\bm{f}}^{(i)})$. the encoded $i^{th}$ frame.
${\bm{z}}^{(0)}$ Section 3.1, Equation 1 encoded first frame ${\bm{f}}^{(0)}$
${\bm{z}}_{\text{pad}}^{(0)}$ Equation 1, Figure 2 padded ${\bm{z}}^{(0)}$ by repeating itself along the first dimension N times.
$t$ Section 3.1, Equation 1 level used by the EDM noise scheduler do determine the noise level
$\hat{{\bm{z}}}_{t}$ Section 3.1, Equation 1 latent representation of frames corrupted by noise level $t$
${\bm{c}}^{(0)}$ Section 3.1, Equation 1 CLIP encoding of the first frame ${\bm{f}}^{(0)}$
${\bm{f}}_{bbox}$ Figure 2 bounding box frames. Each frame contain pixel renderings of bounding boxes on a blank background. Used as conditioning for Box2Video
${\bm{b}}$ Figure 2 latent representation of ${\bm{f}}_{bbox}$ (encoded by $\mathcal{E}$)
${\bm{b}}^{(0)}$ Figure 2 latent representation of the first bounding box frame
${\bm{b}}^{(N-1)}$ Figure 2 latent representation of the last bounding box frame
$\hat{{\bm{b}}}_{t}$ Figure 2 ${\bm{b}}$ corrupted by noise level $t$
Table 4: Glossary of terms.
Appendix C Implementation Details
C.1 Step-by-Step Guide to Plotting bounding box Frames
We create bounding box plots for each frame according to the following steps:
@@ C.1 Step-by-Step Guide to Plotting bounding box Frames
    For each object, we fill the 2D bounding box region with its track ID’s color at 75% transparency. This transparency allows the overlapping regions of the boxes to remain visible.
@@
    If 3D bounding box information is available for an object, we draw the 3D bounding boxes and outline them with the color corresponding to the object’s class label.
@@
    If 3D bounding box information is unavailable, we outline the 2D bounding box with the color corresponding to the object’s class label. Moreover, we mark an X at the rear face of each 3D bounding box for enhanced contextualization.
@@
    In certain experiments, we substitute bounding box plots with trajectory plots. To create a trajectory plot, we first compute the mid-points of its 2D bounding boxes. Subsequently, at each bounding box midpoint, we draw a circle with a diameter of 10 pixels, filled with the color corresponding to its track ID. Within this circle, we draw a smaller circle with a diameter of 5 pixels, filled with the color representing the object’s class label.
C.2 bounding box Visualization and Reconstruction Analysis
In this work, we refrain from introducing new modules dedicated to encoding the bounding box frames. Instead, we leverage the encoder ($\mathcal{E}$) within the SVD’s autoencoder framework. Our experiments demonstrate the efficacy of the SVD autoencoder in accurately encoding and decoding bounding box frames.
@@
Figure 6: Visualizing a BDD100K 2D-bbox frame created using the method described in C.1
@@
Figure 7: Visualizing a VKITTI 3D-bbox frame created using the method described in C.1
Figure 8: Visualizing a BDD100K trajectory frame created using the method described in C.1.
C.3 Dataset Configuration
KITTI: The downloadable content of the official KITTI dataset does not include bounding box label files for the testing set. As a result, we manually partition the dataset’s training folder into two segments for our experiments. The training split encompasses samples ‘$0$’ to ‘0018’, while the testing split consists of samples ‘0019’ and ‘0020’.
@@ C.3 Dataset Configuration
nuScenes: We adhere to the official train/validation split (700/150 scenes) for our training and testing phases.
In our experiments, we use a standardized frame size of $312\times 520$ pixels and a clip length of 25 frames at a frame rate of 7 frames per second. All experiments are trained using the training set and assessed using the validation or test sets. Unless otherwise stated, the results and the visualizations provided are derived from the test sets.
C.4 Training Configuration
Our pretrained SVD models are adaptations from HuggingFace’s diffuser library (von Platen et al., 2022). Specifically, we utilize the ‘stable-video-diffusion-img2vid-xt’ SVD variant in this project.
@@ C.4 Training Configuration
The 3D bounding box generator anticipates frames using 3D bounding boxes, while the 2D generator does the same for 2D bounding boxes. A 1-to-1 model generates frames between the first and last bounding box frames, while a 3-to-1 model generates frames between the first 3 and last bounding box frames. Additionally, there’s a bounding box-trajectory model where the last conditional frame is a trajectory frame instead of a bounding box frame.
@@
For the bounding box generators trained on KITTI or vKITTI datasets, each model undergoes 5 epochs of training. The 2D bounding box generator (1-to-1) is trained on BDD100K for 33,000 iterations. Additionally, the 2D bounding box (3-to-1) and 2D bounding box-trajectory (3-to-1) models start their fine-tuning process for an additional 15,000 iterations from where the 2D bounding box generator (1-to-1)’s training ends. All of these generation models are trained on a single A100-80G GPU with a batch size of 1. We also set gradient accumulation steps to 5 and employed AdamW (Loshchilov & Hutter, 2019) as our optimizer.
@@
The fine-tuning and training procedures for the baseline Stable Diffusion and the ControlNet models are almost the same as of the bounding box generator model. The only difference is that we trained our baseline and ControlNet for 48,000 iterations on BDD100K.
C.5 Model Capacity
The number of parameters for each model, along with its submodule parameter counts, is provided in Table 5.
Model Submodule Status umber of Parameters
BBox Generator VAE-Encoder Frozen 34,163,592
VAE-Decoder Frozen 63,579,183
CLIP-Image Encoder Frozen 632,076,800
ST Condition UNet Trainable 1,524,623,082
Box2Video VAE-Encoder Frozen 34,163,592
VAE-Decoder Frozen 63,579,183
CLIP-Image Encoder Frozen 632,076,800
ST Condition UNet Frozen 1,524,623,082
ControlNet Trainable 680,946,897
Table 5: This table presents the parameter counts for the BBox Generator (2.25 billion) and Box2Video (2.94 billion) models, with a detailed submodule parameter breakdown.
Appendix D Evaluation Details
D.1 Generation Quality Metrics: PSNR, SSIM, LPIPS and FVD
@@
PSNR, or Peak signal-to-noise ratio, calculates the ratio between the maximum value of a signal and the noise that affects the fidelity of its representation. It is closely related to Mean Squared Error (MSE) loss. It can be in fact calculated by subtracting the log of MSE from a constant. It is long been used as a metric for image reconstruction quality for its ease of use and mathematical properties. However, PSNR suffers multiple issues as a metric for image generation tasks. Its insensitiveness to various distortion such as blurring, pepper noise, and mean-shifting means that image pairs that look very different to human can have very similar PSNR scores (Wang & Bovik, 2009; Wang et al., 2004a).
@@
SSIM, or Structural similarity index measure, is a metric that aims to evaluate similarity in ”structures” of image pairs – as opposed to absolute errors as in the cases of MSE and PSNR. Because the human visual system is more sensitive to structural distortions, this metric is more closely aligned with human evaluation compared to PSNR. However, some visually apparent distortions are not captured by the SSIM. Things like changing hue and brightness do not seem to affect SSIM much (Kotevski & Mitrevski, 2010) and Sharif et al. (2018) presents more examples of unpredictable ssim behaviours when presented with different distortions.
@@
LPIPS, or Learned Perceptual Image Patch Similarity (Zhang et al., 2018), is a neural network based approach to accessing image similiarity. It does so by computing distance between some reprensentation of image patches. These representations are obtained by running image patches through some pre-defined neural network. LPIPS has been shown to math human perception well.
@@
FVD, or Frechet Video Distance (Unterthiner et al., 2019), is a metric for generative video models. It is based on the Fréchet inception distance (FID) for images. Unlike the 3 other metrics that focuses on image frames and do not consider their temporal relationships, FVD is specifically designed for videos. It takes into consideration the visual quality, temporal coherence, and diversity of samples of videos generated by the model under evaluation. The large scale human study carried out by the authors suggest FVD consistently agrees with human evaluation of videos.
D.2 Prediction Scores: Converting bounding box Frames to Binary Masks
@@
To transform generated bounding box frames into binary masks, we first need to perform post-processing on the frames. These frames originate from a network output, and although background pixels that seem black, they may not all be precisely zero. To prevent these pixels being detected as bounding box masks, we aggregate the pixel values across color channels, and for those channels whose sum is below 50, we convert them to black pixels. Subsequently, we convert the frames into binary masks by aggregating all channels and changing the colored pixel values to 1.
@@
Accurately evaluating bounding box generations remains challenging. We recognize that our current metrics (maskIoU, maskR and maskP), which rely on binary masks to evaluate the performance of bounding box generators, have limitations as they do not consider object tracking IDs (the colors of the bounding boxes). However, we still believe that the metric evaluations using binary masks can offer valuable insights into our model’s performance.
D.3 Motion Control Assessment: YOLOv8 Object Detector Configurations
@@
We utilize the “yolov8x” variant of the YOLOv8 model, implemented by Ultralytics (Jocher et al., 2023), for object detection. We set the detector’s Non-Maximum Suppression Intersection Over Union (NMS-IoU) threshold to be 0.35 across all experiments.
Since YOLOv8 was trained to detect objects from the MS COCO dataset (Lin et al., 2015), we adjust the class labels of our dataset to match the MS COCO labels using the following mappings:
@@
nuScenes human.{adult, child, construction_worker, personal_mobility, police_officer, wheelchair}–person; vehicle.{bicycle, motorcycle}–person; vehicle.{bus, construction, ambulance, police, trailer, truck}–truck; car–car
@@
We have attempted to enhance our detection results across frames by using a tracker in conjunction with the YOLOv8 detector. Specifically, we use the BoT-SORT (Aharon et al., 2022) tracker. However, we’ve observed an increase in the precision but a significant drop in the recall of the detection scores. Thus, we’ve reverted to using only the YOLOv8 detection model to compute the AP metrics.
D.4 Motion Control Assessment: Average Precision Calculation
@@
We begin by employing a YOLOv8 detection model on both the generated and ground-truth videos, yielding a series of bounding box detections for each. When we generate bounding box detections on ground-truth videos, we keep the confidence cutoff fixed at 0.6.
Next, we calculate the AP scores at different IoU thresholds following the MS COCO protocol (Lin et al., 2015): compute 10 AP scores at IoU thresholds ranging from 0.5 to 0.95, in increments of 0.05. The AP score computed at threshold $\tau$ is referred to as $\text{AP}_{\tau\times 100}$. Mean average precision (mAP) is the average of all the AP scores. The computed results are reported in Table 4.4.
@@
Appendix E Generation Results
E.1 Stable Video Diffusion vs Box2Video: Generation Quality Comparison
The following are generation comparisons between the fine-tuned Stable Video Diffusion baseline model and the Box2Video conditioned on the ground-truth bounding box frames. All visualizations are generated using the BDD100k dataset.